Consider a population consisting of 2.3.4 and 6 illustrate the situation by doing the following:





1.Compute the mean and variance of the population





2.list all possible samples of size 2





3.construct the sampling distribution of the sample





4. Compute the mean and standard deviation of the sampling distribution of the sample means





5.construct the histogram.



1
Expert's answer
2022-05-17T07:57:30-0400

We have population values 2,3,4,6, population size N=4 and sample size n=2.

1. Mean of population "(\mu)" = "\dfrac{2+3+4+6}{4}=3.75"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{4}(3.0625+0.5625"


"+0.0625+5.0625)=2.1875=\dfrac{35}{16}"

"\sigma=\sqrt{\sigma^2}=\sqrt{2.1875}\approx1.47902"

2. Select a random sample of size 2 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{4}C_2=6."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 2,3 & 5/2 \\
 \hdashline
 2 & 2,4 & 6/2 \\
 \hdashline
 3 & 2,6 & 8/2 \\
 \hdashline
 4 & 3,4 & 7/2 \\
 \hdashline
 5 & 3,6 & 9/2 \\
 \hdashline
 6 & 4,6 & 10/2 \\
 \hdashline
\end{array}"



3.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 5/2 & 1/6 & 5/12 & 25/24 \\
 \hdashline
 6/2 & 1/6 & 6/12 & 36/24 \\
 \hdashline
 7/2 & 1/6 & 7/12 & 49/24 \\
 \hdashline
 8/2 & 1/6 & 8/12 & 64/24 \\
 \hdashline
 9/2 & 1/6 & 9/12 & 81/24 \\
 \hdashline
 10/2 & 1/6 & 10/12 & 100/24 \\
 \hdashline
\end{array}"



4. Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{45}{12}=\dfrac{15}{4}=3.75=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{355}{24}-(\dfrac{15}{4})^2=\dfrac{35}{48}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{35}{48}}\approx0.85391"

5.